OpenStudy (anonymous):

At what height does a 1000-kg mass have a potential energy of 1 J relative to the ground? Show your work.

OpenStudy (raffle_snaffle):

Are we trying to find the gravitational work on the 1000kg mass?

OpenStudy (anonymous):

Tbh I have no idea whatsoever >.< lol

OpenStudy (anonymous):

Potential Energy is:\[E_p=m·g·h \rightarrow h=\frac{ E_p }{ m·g }=\frac{ 1 }{ 1000·9.81 }=0.0001m\]and 0.0001m=0.1 mm

OpenStudy (anonymous):

I don't understand what I'm supposed to plug into that eaquation

OpenStudy (festinger):

It really depends on how you want to calculate it. For most cases we use newton's law of universal attraction to calculate the work, but in this case Gravitational Energy = mass * gravitational constant * height will do. This question wants us to find height. We have mass = 1000kg, E = 1J, and g = 9.81m/s^2, since we are dealing with on earth. Putting it all together should give you the solution stated by CarlosGP

OpenStudy (anonymous):

\[1_{p} = 1000kg \times 9.81m \div s ^{2} \times h \rightarrow h = \frac{ 1 _{p} }{1000kg \times 9.81m/s ^{2} }\] Is that even close to a start

OpenStudy (festinger):

Yes, but there is no subscript p, but there should the the unit of Joule, which is actually: kg (m/s)^2, putting it all together h has the unit of m as expected.

OpenStudy (anonymous):

Can I please see the correct order please? I'm slightly confused with the second part of what you just said.

OpenStudy (festinger):

Maybe in a few days, I tried typing it out in TeX twice but it crashed, taking everything I typed with it.

OpenStudy (anonymous):

Okay then... thank you for trying T:

OpenStudy (anonymous):

@AngelCriner Go back to the formula I wrote: Ep=Potential Energy in Joules h=height in meters m=mass in Kilograms g=9.81 m/sec^2

OpenStudy (anonymous):

\[1J=1000kg \times 9.81m/\sec^2 \times ? \rightarrow ? = \frac{ 1J }{ 1000 \times 9.81 } = \frac{ 1}{ 1000 \times 9.81 } = 0.0001m\] I am extremely lost

OpenStudy (anonymous):

? = h = height in meters

OpenStudy (anonymous):

I didn't know where the height came from though so I didn't have a number to plug in >.<

OpenStudy (festinger):

Work is force x distance, and the unit is joule (energy) force is mass x acceleration, and unit is newton (force) acceleration has units of m/s^2 square brackets means units of: \[ [Energy] = J = [Force]*[Distance] = [mass]*[accleration]*[dist] = kg * m / s^2 * m\] which simplifies to: \[ J = kg * m^2/s^2\] Substituting this J into the original expression you get height in units of meters, as it should be.

OpenStudy (anonymous):

Read the problem and note that height is precisely the unknown you need to calculate

OpenStudy (anonymous):

Oh.My.Goodness. >_\ Oh...my...GOODNESS. I'm so sorry >.< I'm failing to understand the correlation between the problem and the expression. /: